OpenStudy (anonymous):

Becky has $22, and she earns $8 an hour at her new job. The equation m = 8h + 22 can be used to determine how much money in dollars [m] Becky has after any number of hours at work [h]. After how many hours of work will Becky have $78? Select one: 6 8 7 5 Solve each system of equations using the method of your choice: Substitution, Elimination, or Graphing. Y = x + 4 y = 4x + 1 Select one: (5, 3) (-1, 3) (1, 5) (-5, -1) Solve each system of equations using the method of your choice: Substitution, Elimination, or Graphing. y = x + 1 2x + y = 10 Select one: (1, 5) (3,

OpenStudy (anonymous):

To solve number 1, do 78=8h+22 subtract 22 56=8h Divide by 8 h=7 Before I go on, did you understand that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Alright, I wrote that down so I have it

OpenStudy (anonymous):

Ok, for two, let's graph. The answer is where the lines intersect, which is at (1,5).

OpenStudy (anonymous):

Anyways, for three, graphing it shows the answer is (3,4)

OpenStudy (anonymous):

OK @zeinas.

OpenStudy (anonymous):

If asdkfj...doesn't care, please get off the site. This is an abuse.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

go get life @asdijhflakjsdh

OpenStudy (anonymous):

On the second problem, both are in the form of y=something. Put both equations equal to each other and solve for x. Once you have x, you can get y by solving.

OpenStudy (anonymous):

Ok I'll go that @dss

OpenStudy (anonymous):

*do and I got (1,5)

OpenStudy (anonymous):

asdij...I may be old, but I can do math in my sleep. Obviously, I'm a lot smarter than you. And have a larger vocabulary. Get off the site and don't waste our efforts.

OpenStudy (anonymous):

yes, (1,5) is right.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@Yolo4mecuite do you have any other questions, if not can you close this one so we don't have deal with that stupid guy?

OpenStudy (anonymous):

y = x + 1 2x + y = 10 with this problem, I'd use substitution. Substitute the 1st equation into the second. You'll have: 2x+(x+1)=10 Now solve for x

OpenStudy (anonymous):

Well I actually do but I'll make another post on it. and ok @ds

OpenStudy (anonymous):

@dss

OpenStudy (anonymous):

ok